OpenStudy (anonymous):

find the inverse of the one-to-one function f(x)=4/5x-1

OpenStudy (anonymous):

f-1(x)=4/5x+1/5 is answer on the key

OpenStudy (welshfella):

find x in terms of f(x) then replace f(x) by x

OpenStudy (anonymous):

y=4/5x-1

OpenStudy (anonymous):

x=4/5y-1

OpenStudy (anonymous):

now im lost

OpenStudy (welshfella):

is the right side